### v2

@@ 6 Numerical Experiments
Figure 2: Prediction accuracy of nonparametric S-RAM with underlying parametric S-RAM data. Blue dots are ground truths. Red ranges are prediction intervals for the predicted flow using nonparametric S-RAM in path problems. Red dots are optimistic estimates of total welfare using nonparametric S-RAM in assignment problems.
7 Conclusion and Future Work
@@
In this paper, we develop a nonparametric framework for counterfactual prediction in combinatorial problems based on a separable representative agent model (S-RAM). We establish an exact characterization of choice probabilities representable under S-RAM over general 010-1 polytopes and show that consistency verification reduces to solving a polynomial-sized linear program. Building on this, we propose a nonparametric tractable method for counterfactual prediction. For S-RAM-inconsistent data, we introduce a compact mixed-integer convex formulation to compute the best-fitting S-RAM approximation. We conduct extensive numerical experiments that demonstrate the strong empirical performance of the proposed method. Finally, we highlight several challenges that underscore future research directions. Developing advanced methods for handling large scale instances where the proposed integer programs formulations may fail to scale well is an important direction for future research. Another direction is to extend the framework to settings where a compact 011 polytope representation is not available, such as the Traveling Salesman Problem (TSP), or even beyond 011 polytopes, as in the knapsack problem, thereby accommodating richer combinatorial structures, and broadening the framework’s applicability.
@@ Appendix A Missing Proofs of the Paper
$\nu_{j}-c_{j}^{\prime}(p_{j}^{k})+\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}+\mu_{j}^{k}-w_{j}^{k}=0,\quad\forall j\in\mathcal{I}^{(k)},$ (11)
$\mu_{j}^{k}x_{j}^{*}=0,\ w_{j}^{k}(1-x_{j}^{*})=0,\quad\forall j\in\mathcal{I}^{(k)}.$ (12)
@@
Figure 3: An illustration of the construction of the functions $c_{j}^{\prime}$ when: (a) $p_{j}^{K_{j}}<1$ and (b) $p_{j}^{1}>0$.
@@
This is possible in polynomial time by solving a linear program where the above conditions are formulated as constraints and maximizing $\epsilon.$ This linear program involves $mK$ variables for $(\lambda_{i}^{k}\in\mathbb{R}:i\in[m],k\in\mathcal{K})$ and one variable for $\epsilon,$ and $\mathcal{O}((n+m)K)$ constraints.
{proofof}
@@
Therefore, the optimal value of the formulations in Propositions 4.1 and 4.2 are identical.
{proofof}
@@
Therefore, the optimal value of the formulations in Propositions 5.1 and 5.2 are identical.
Appendix B Details on Numerical Experiments in Section 5
B.1 Data Generation
Gaussian Data for representability and estimation test. We provide a detailed description of the data generation procedure used in our experiments for evaluating the representational power of S-RAM across three combinatorial problems: longest path, shortest path, and assignment.
@@ B.1 Data Generation
For an instance $\boldsymbol{p}^{\mathcal{K}}$, we verify whether the observed data $\boldsymbol{p}^{\mathcal{K}}$ is representable by S-RAM by checking the conditions in equation (3). If the condition is violated, we compute the average absolute deviation loss using equation (9), where the loss function is defined as $\sum_{k\in\mathcal{K}}\sum_{e\in\boldsymbol{E}^{k}}|x_{e}-p_{e}|.$ Here, $p_{e}$ is the observed flow on edge $e$, and $x_{e}$ is the flow estimated under the S-RAM model. Finally, we report, for each $K$ and noise setting, the proportion of representable instances and the average absolute deviation loss across the 20 generated instances.
@@
For the path problems, we consider both the longest path and shortest path formulations on directed acyclic graphs with $n=10$ nodes. Graphs are constructed in a similar way as in earlier experiments. To ensure connectivity from source to sink, we first create a fixed path from node 0 (the source node) to node n1n-1 (the sink node) and include all edges along this path. Additional edges are then included independently with probability sampled uniformly from the interval $[0.4,0.6]$. Edge costs are modeled as the sum of a deterministic term and a noise term. To ensure that a specific edge of interest consistently appears in the optimal solution, we set its deterministic cost to 8 in the longest path problem and to 1 in the shortest path problem. For all other edges, the deterministic costs are sampled independently from the uniform distribution over $[1,7]$. The noise on each edge follows an exponential distribution, with nonidentical rates sampled from the interval $[1,2]$. This results in parametric MDM data with nonidentical exponential marginal distributions.
@@
For each value of $K$, we generate $K$ historical graphs $\boldsymbol{G}^{1}=(\boldsymbol{V}^{1},\boldsymbol{E}^{1}),\boldsymbol{G}^{2}=(\boldsymbol{V}^{2},\boldsymbol{E}^{2}),\cdots,\boldsymbol{G}^{K}=(\boldsymbol{V}^{K},\boldsymbol{E}^{K})$, following the procedure above. For a graph $\boldsymbol{G}^{k}$ with $k\in\mathcal{K}$, we solve a convex program defined by the corresponding parametric MDM model to obtain the choice probabilities. Let $\lambda_{ij}$ be the rate parameters for each edge $(i,j)$ in the set of all edges $\boldsymbol{E}$. From (Natarajan et al., 2009), we have that, for a graph $\boldsymbol{G}^{k}$, the convex program with nonidentical exponential marginal distributions for the longest path problem is as follows:
$\max_{\boldsymbol{x}}$ $\sum_{(i,j)\in\boldsymbol{E}^{k}}(\nu_{ij}+\frac{1}{\lambda_{ij}})x_{ij}-\sum_{(i,j)\in\boldsymbol{E}^{k}}\frac{1}{\lambda_{ij}}x_{ij}\ln{x_{ij}}$
s.t. $\sum_{j:(i,j)\in\boldsymbol{E}^{k}}x_{ij}-\sum_{j:(j,i)\in\boldsymbol{E}^{k}}x_{ji}=\begin{cases}1,\quad&i=s,\\ -1,\quad&i=t,\\ 0,\quad&i\in\boldsymbol{V}^{k},\end{cases}$
$x_{ij}\geq 0\quad\forall\ (i,j)\in\boldsymbol{E}^{k}.$
From (Natarajan et al., 2009), we have that, for a graph $\boldsymbol{G}^{k}$ with $k\in\mathcal{K}$, the convex program with nonidentical exponential marginal distributions for the shortest path problem is as follows:
$\min_{\boldsymbol{x}}$ $\sum_{(i,j)\in\boldsymbol{E}^{k}}(\nu_{ij}+\frac{1}{\lambda_{ij}})x_{ij}+\sum_{(i,j)\in\boldsymbol{E}^{k}}\frac{1}{\lambda_{ij}}(1-x_{ij})\ln{(1-x_{ij})}$
s.t. $\sum_{j:(i,j)\in\boldsymbol{E}^{k}}x_{ij}-\sum_{j:(j,i)\in\boldsymbol{E}^{k}}x_{ji}=\begin{cases}1,\quad&i=s,\\ -1,\quad&i=t,\\ 0,\quad&i\in\boldsymbol{V}^{k},\end{cases}$
$x_{ij}\geq 0\quad\forall\ (i,j)\in\boldsymbol{E}^{k}.$
These choice probabilities are used as the observed data $\boldsymbol{p}^{\mathcal{K}}$ to train the nonparametric S-RAM model. For prediction, we generate a new graph $\boldsymbol{G}^{\text{new}}$ and evaluate whether the model can predict the expected flow on the pre-specified key edge, which remains influential in the new graph.
For the assignment problem, we generate historical graphs by starting from a complete bipartite graph with 4 nodes on the left and 6 nodes on the right, and then randomly designating a subset of nodes as unavailable to generate $K$ historical graphs $\boldsymbol{G}^{1}=(\boldsymbol{V}^{1},\boldsymbol{E}^{1}),\boldsymbol{G}^{2}=(\boldsymbol{V}^{2},\boldsymbol{E}^{2}),\cdots,\boldsymbol{G}^{K}=(\boldsymbol{V}^{K},\boldsymbol{E}^{K})$, for each value of $K$. For each historical graph, we remove all edges associated with the unavailable nodes. This yields a sequence of structurally varying graphs over which we compute observed data $\boldsymbol{p}^{\mathcal{K}}$. The prediction task is then to estimate the total system welfare on a new graph $\boldsymbol{G}^{\text{new}}$ where node availability is unseen from the historical graphs.
Edge costs in the assignment setting follow a parametric MDM structure. Each deterministic edge cost $v_{e}$ is sampled independently from the interval $[1,7]$. The noise term $\epsilon_{e}$ is sampled from a uniform distribution with nonidentical support. For each edge, the lower bound $a$ is sampled from $[1,2]$ and the upper bound $b$ is sampled from $[5,10]$, so that $\epsilon_{e}\sim\text{Uniform}(a,b)$. This construction results in nonidentical uniform marginal distributions across edges while preserving the MDM structure.
For each historical graph $\boldsymbol{G}^{k}$, we solve the corresponding convex program induced by the parametric MDM to obtain the choice probabilities $\boldsymbol{p}^{\mathcal{K}}$. From (Natarajan et al., 2009), we have that, for a graph $k$, the convex program to compute the choice programs with nonidentical uniform marginal distributions for the assignment is as follows:
$\max_{\boldsymbol{x}}$ $\sum_{i\in\boldsymbol{V}_{L}^{k}}\sum_{j\in\boldsymbol{V}_{R}^{k}}(\nu_{ij}+b_{ij})x_{ij}+\sum_{i\in\boldsymbol{V}_{L}^{k}}\sum_{j\in\boldsymbol{V}_{R}^{k}}\frac{b_{ij}-a_{ij}}{2}x_{ij}^{2}$
s.t. $\sum_{j\in\boldsymbol{V}_{R}^{k}}x_{ij}=1,\quad\forall i\in\boldsymbol{V}_{L}^{k},$
$\sum_{i\in\boldsymbol{V}_{L}^{k}}x_{ij}=1,\quad\forall j\in\boldsymbol{V}_{R}^{k},$
$x_{ij}\geq 0\quad\forall\ i\in\boldsymbol{V}_{L}^{k},j\in\boldsymbol{V}_{R}^{k}.$
These are used as observed data $\boldsymbol{p}^{\mathcal{K}}$ for training the nonparametric S-RAM model. For prediction, we evaluate whether the nonparametric S-RAM can accurately predict the total expected system welfare, defined as the same objective above with the new graph $\boldsymbol{G}^{\text{new}}$ in the selected matching.
B.2 Representability, Estimation, Prediction Formulations
B.2.1 Longest Path Problem
With a slight abuse of notation, suppose that $\boldsymbol{G}=(s\cup t\cup\boldsymbol{V},\boldsymbol{E})$ is the complete DAG that describes all possible nodes and edges of the network, and we have flow data for a collection $K$ DAGs, denoted as $\boldsymbol{G}^{1},\boldsymbol{G}^{2},\cdots,\boldsymbol{G}^{K},$ with $\boldsymbol{G}^{k}=(s\cup t\cup\boldsymbol{V}^{k},\boldsymbol{E}^{k})$, for every $k\in\mathcal{K}$. For a graph $\boldsymbol{G}^{k}$, let $p_{ij}^{k}\in[0,1]$ denote the flow of edge $(i,j)$ in $\boldsymbol{G}^{k}$. Given observed flow data $\boldsymbol{p}^{\mathcal{K}}=(p_{ij}^{k}:(i,j)\in\boldsymbol{E}^{k},k\in\mathcal{K})$, by Theorem 3.1, the S-RAM characterization for longest path problems is given below.
S-RAM characterization for Longest Path Problems. Under Assumption 1, an observed flow collection $\boldsymbol{p}^{\mathcal{K}}$ is representable by an S-RAM if and only if there exists $(\lambda_{i}^{k}\in\mathbb{R}:i\in s\cup t\cup\boldsymbol{V}^{k},k\in\mathcal{K})$, such that for any two distinct graph $k,l\in\mathcal{K}$ containing a common edge $(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},$
$\begin{aligned} &\lambda_{j}^{k}-\lambda_{i}^{k}>\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad p_{ij}^{k}>p_{ij}^{l},\\ &\lambda_{j}^{k}-\lambda_{i}^{k}=\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad 0<p_{ij}^{k}=p_{ij}^{l}<1.\end{aligned}$ (13)
@@ B.2.1 Longest Path Problem
$\begin{aligned} \max_{\epsilon,\boldsymbol{\lambda}}\quad&\epsilon\\ \text{s.t.}\quad&\lambda_{j}^{k}-\lambda_{i}^{k}\geq\lambda_{j}^{l}-\lambda_{i}^{l}+\epsilon\quad\text{if}\quad p_{ij}^{k}>p_{ij}^{l},\quad\forall k,l\in\mathcal{K},\forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\lambda_{j}^{k}-\lambda_{i}^{k}=\lambda_{j}^{l}-\lambda_{i}^{l}\quad\text{if}\quad 0<p_{ij}^{k}=p_{ij}^{l}<1,\quad\forall k,l\in\mathcal{K},\forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l}.\end{aligned}$ (14)
Predictions for Longest Path Problems. Suppose that $\boldsymbol{p}^{\mathcal{K}}$ is representable by S-RAM, given a new graph $\boldsymbol{G}^{\text{new}}=(s\cup t\cup\boldsymbol{V}^{\text{new}},\boldsymbol{E}^{\text{new}})$, by Proposition 4.2, the prediction formulation to compute $\underline{f}_{\text{new}}$ for the longest path problem by taking $0<\epsilon<1/(2|\boldsymbol{V}|K),$ is as follows.
$\min_{\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{\delta}\!\!\,}\quad$ $\quad f(\boldsymbol{x})$ (15)
Likewise, the best-case estimate for the objective value $\bar{f}_{\text{new}}$ equals the optimal value obtained by maximizing over the constraints in the above mixed integer linear program. In our experiment, we take $f(\boldsymbol{x})=x_{\hat{e}}$ where $\hat{e}$ is the pre-specified edge.
Estimation for Longest Path Problems. In the case of longest path problems, by taking $0<\epsilon<1/(2|\boldsymbol{V}|K),$ formulation (9) can be written as:

### v3

@@ 6.2.2. Predictive ability of s-RAM in the shortest path problem
Figure 3. Prediction accuracy of nonparametric s-RAM on data generated from an underlying parametric s-RAM. Blue dots denote ground-truth values. Red dots indicate optimistic estimates of bottleneck congestion in the shortest path problem by using nonparametric s-RAM.
6.3. Assignment Problem
Next, we evaluate the performance of s-RAM for the shortest path problem. Due to space limitations, we refer the reader to Section B.2.2 for details on the representability check, estimation, and prediction formulations, and to Section B.1 for a complete description of the data generation.
6.3.1. Explanatory ability of s-RAM in the assignment problem
Generation of historical $\mathcal{X}^{(k)}$ instances. For the assignment problem, we consider bipartite graphs with $4$ nodes on the left and $6$ nodes on the right. The number of historical instances varies across $K\in\{5,10,15,20\}$. Edges are included independently with probabilities drawn uniformly from $[0.4,0.6]$, subject to the constraint that each node is incident to at least one edge, ensuring feasibility of the assignment problem. On average, the resulting graphs contain approximately $12$ edges.
Generation of historical $\boldsymbol{p}^{(k)}$ instances. The generation of $\boldsymbol{p}^{(k)}$ for the assignment problem follows the same procedure as in the longest path experiments, with edge costs given by $v_{e}+\epsilon_{e}$ under the same noise structures and parameter settings. The only difference is that the deterministic assignment problem is solved to obtain the observed data.
Representability test and estimation with $\boldsymbol{p}^{\mathcal{K}}$. The results for the assignment problem are reported in Table 4. Consistent with the longest and shortest path experiments, s-RAM exhibits strong explanatory ability across all tested scenarios despite model misspecification. In particular, the fraction of s-RAM representable instances remains high throughout the tested range of $K$. For example, when $K=5$, at least $95\%$ of the instances are s-RAM representable under all noise structures, and representability remains above $60\%$ even when $K=20$. Moreover, the associated estimation error is extremely small: the average absolute deviation loss is consistently on the order of $10^{-5}$ or smaller across all settings, including cases where exact representability is not achieved. As in the path problems, representability decreases with $K$ due to the growing number of feasibility constraints induced by additional historical observations. Nevertheless, the assignment problem appears particularly well-suited to the s-RAM framework, with robust explanatory performance and negligible estimation error across a wide range of correlation structures.
Table 4. Representation and estimation in the assignment problem with Gaussian instances
K independent gaussian gaussian with -ve correlation gaussian with +ve correlation
fraction
of s-RAM
representable
instances
average absolute
deviation loss (se)
$(10^{-5})$
fraction
of s-RAM
representable
instances
average absolute
deviation loss (se)
$(10^{-5})$
fraction
of s-RAM
representable
instances
average absolute
deviation loss (se)
$(10^{-5})$
5 0.95 0.50 (0.49) 1.00 0.00 (0.00) 0.95 0.10 (0.10)
10 0.90 1.55 (1.24) 0.95 1.40 (1.36) 0.85 3.15 (2.92)
15 0.85 4.03 (2.45) 0.85 2.50 (1.82) 0.90 0.20 (0.14)
20 0.75 2.38 (1.75) 0.60 8.53 (4.40) 0.60 1.97 (1.13)
6.3.2. Predictive ability of s-RAM in the assignment problem
We evaluate the predictive accuracy of the nonparametric s-RAM on counterfactual tasks for the assignment problem using data generated from parametric s-RAMs. In this setting, we examine whether s-RAM can accurately predict the total system welfare on a new graph when some nodes become unavailable.
Generation of training instances $\boldsymbol{p}^{(k)}$ and ground truth in the assignment problem. In assignment, we let the noise marginals be non-identical and uniformly distributed, with lower bounds drawn from $[1,2]$ and upper bounds drawn from $[5,10]$. Counterfactual instances are generated by randomly designating a subset of nodes as unavailable in the bipartite graph and removing all edges incident to those nodes. Then we use formulation (20) to generate training instances $\boldsymbol{p}^{(k)}$. Specifically, let $\boldsymbol{\nu}$ denote the deterministic weight matrix, and let $\mathbf{a}$ and $\mathbf{b}$ denote the matrices of lower and upper bounds of the uniform noise distributions, respectively. Given a new bipartite graph $\boldsymbol{G}^{\text{new}}=(\boldsymbol{V}_{L}^{\text{new}},\boldsymbol{V}_{R}^{\text{new}},\boldsymbol{E}^{\text{new}})$, let $\boldsymbol{x}^{\text{true}}$ denote the ground-truth matching rates for $\boldsymbol{G}^{\text{new}}$. We then compute the ground-truth system welfare for each counterfactual instance with formulation (20) under the same parameter values as in the training instances. The corresponding ground-truth system welfare is given by
$\sum_{i\in\boldsymbol{V}_{L}^{\text{new}}}\sum_{j\in\boldsymbol{V}_{R}^{\text{new}}}(\nu_{ij}+b_{ij})x_{ij}^{\text{true}}-\sum_{i\in\boldsymbol{V}_{L}^{\text{new}}}\sum_{j\in\boldsymbol{V}_{R}^{\text{new}}}\frac{b_{ij}-a_{ij}}{2}{(x_{ij}^{\text{true}})}^{2}.$
Prediction with nonparametric s-RAM in the assignment problem. For the assignment problem, computing the worst-case welfare prediction leads to a nonconvex mixed-integer optimization problem, which is computationally challenging and may be unreliable when solved using standard solvers. We therefore focus on the best-case s-RAM estimate for the total system welfare, defined as the maximum achievable welfare over all s-RAMs consistent with the observed data and the counterfactual graph structure. Formally, this quantity is given by
$\max_{\boldsymbol{x}\in\mathcal{U}{\text{new}}}\sum_{i\in\boldsymbol{V}_{L}^{\text{new}}}\sum_{j\in\boldsymbol{V}_{R}^{\text{new}}}(\nu_{ij}+b_{ij})x_{ij}-\sum_{i\in\boldsymbol{V}_{L}^{\text{new}}}\sum_{j\in\boldsymbol{V}_{R}^{\text{new}}}\frac{b_{ij}-a_{ij}}{2}x_{ij}^{2}.$
Recall that $\mathcal{U}{\text{new}}$ denotes the feasible s-RAM nonparametric prediction region defined in Section 4.1.
As shown in Figure 4, across all test instances, the best-case welfare predicted by s-RAM closely matches the ground-truth system welfare, indicating strong predictive performance of the nonparametric approach in the assignment setting.
@@ 6.3.2. Predictive ability of s-RAM in the assignment problem
Figure 4. Prediction accuracy of nonparametric s-RAM on data generated from an underlying parametric s-RAM. Blue dots denote ground-truth values. Red squares indicate optimistic total welfare in the assignment problem by using nonparametric s-RAM.
6.4. Constrained Subset Selection Problem
Next, we evaluate the performance of s-RAM for the shortest path problem. Due to space limitations, we refer the reader to Section B.2.3 for details on the representability check, estimation, and prediction formulations, and to Section B.1 for a complete description of the data generation.
6.4.1. Explanatory ability of s-RAM in the constrained subset selection problem
Generation of historical $\mathcal{X}^{(k)}$ instances. For the constrained subset selection problem, we consider a ground set of available items $\mathcal{N}$ with $|\mathcal{N}|=10$. We focus on a cardinality-constrained setting, where $\mathcal{X}^{(k)}=\left\{\boldsymbol{x}:\sum_{i\in\mathcal{I}^{(k)}}x_{i}\leq C,\ x_{i}\in\{0,1\},\ \forall i\in\mathcal{I}^{(k)}\right\}$ with $C=2$ and $\mathcal{I}^{(k)}\subseteq\mathcal{N}$ for each historical instance $k$. The number of historical instances varies across $K\in\{5,10,15,20\}$. For each instance, we randomly generate the available item set $\mathcal{I}^{(k)}$ with cardinality $|\mathcal{I}^{(k)}|\in\{3,4\}$.
Generation of historical $\boldsymbol{p}^{(k)}$ instances. The generation of $\boldsymbol{p}^{(k)}$ for the constrained subset selection problem follows the same procedure as in the longest path experiments, with item costs given by $v_{i}+\epsilon_{i}$ under the same noise structures and parameter settings as edge costs. The only difference is that the deterministic constrained problem is solved to obtain the observed data.
Representability test and estimation with $\boldsymbol{p}^{\mathcal{K}}$. The results for the constrained subset selection problem are reported in Table 5. Compared with the longest path, shortest path, and assignment problems, s-RAM exhibits weaker explanatory performance in this setting, which is expected since the data are generated outside the s-RAM framework and thus constitute a stress test. In particular, s-RAM is able to represent a nontrivial fraction of instances for small $K$, but representability deteriorates more rapidly as $K$ increases, reflecting the increasing number of feasibility constraints induced by additional historical observations. Nevertheless, the estimation loss remains moderate: the average absolute deviation loss stays on the order of $10^{-2}$ across all settings. This larger scale is natural, as outcomes in the constrained subset selection problem are not normalized to sum to one, unlike the flow- and matching-based problems considered earlier. Overall, despite model misspecification and the absence of normalization, s-RAM continues to provide reasonable explanatory performance in this more challenging setting.
Table 5. Representation and estimation in the constrained subset selection problem with Gaussian instances
K independent gaussian gaussian with -ve correlation gaussian with +ve correlation
fraction
of s-RAM
representable
instances
average absolute
deviation loss (se)
$(10^{-2})$
fraction
of s-RAM
representable
instances
average absolute
deviation loss (se)
$(10^{-2})$
fraction
of s-RAM
representable
instances
average absolute
deviation loss (se)
$(10^{-2})$
5 0.75 0.35 (0.20) 0.90 0.13 (0.09) 0.85 0.25 (0.21)
10 0.50 0.57 (0.20) 0.45 0.73 (0.33) 0.45 0.60 (0.19)
15 0.15 0.66 (0.13) 0.10 0.98 (0.17) 0.20 0.67 (0.17)
20 0.05 1.96 (0.36) 0.10 1.29 (0.27) 0.00 1.31 (0.21)
6.4.2. Predictive ability of s-RAM in the constrained subset selection problem
We evaluate the predictive accuracy of the nonparametric s-RAM on counterfactual tasks for the constrained subset selection problem under a cardinality constraint. In this setting, we examine whether s-RAM can predict the counterfactual marginal allocation of a selected item, defined as its optimal fractional weight in the solution to the constrained subset selection problem on a new instance.
Generation of training instances $\boldsymbol{p}^{(k)}$ and ground truth for the constrained subset selection problem. In the constrained subset selection problem, all items are assigned identical deterministic weights, while the noise marginals follow non-identical exponential distributions with rates sampled uniformly from $(1,2)$. Then, we use formulation (21) to generate training instances $\boldsymbol{p}^{(k)}$. We evaluate the prediction performance of the nonparametric s-RAM across collection sizes $K\in\{20,40,80\}$, generating $20$ training instances for each value of $K$. For a new instance, the ground-truth marginal allocation is obtained by solving the corresponding constrained subset selection problem under the same parameter values with formulation (21). Let $\boldsymbol{x}^{\text{true}}$ denote the resulting vector of marginal allocation of items in a new instance $\mathcal{I}^{(new)}$, and let $i^{*}$ be the item of interest. The corresponding ground-truth quantity is then $x^{\text{true}}_{i^{*}}$.
Prediction with nonparametric s-RAM in the constrained subset selection problem. For the constrained subset selection problem, we compute the s-RAM worst- and best-case estimates for the marginal allocation of the selected feature over all s-RAMs consistent with the observed data. Given a new set $\mathcal{I}^{(new)}$, the worst case estimate is given as $\min_{\boldsymbol{x}\in\mathcal{U}{\text{new}}}x_{i^{*}}$ and the best case estimate is given as $\max_{\boldsymbol{x}\in\mathcal{U}{\text{new}}}x_{i^{*}}$. Recall that $\mathcal{U}{\text{new}}$ denotes the feasible s-RAM nonparametric prediction region defined in Section 4.1.
Figure 5 shows that the estimated marginal allocations closely track the ground truth values across all tested settings. Moreover, the prediction intervals become progressively tighter as the number of historical instances increases, indicating that the nonparametric s-RAM effectively captures the underlying structure of the constrained subset selection problem.
subset selection with $K=20$
subset selection with $K=40$
subset selection with $K=80$
Figure 5. Prediction accuracy of nonparametric s-RAM on data generated from an underlying parametric s-RAM. Blue dots denote ground-truth values. Red ranges represent prediction intervals for the predicted marginal allocation in the constrained subset selection problem by using nonparametric s-RAM.
7. Conclusion and Future Work
@@
In this paper, we develop a nonparametric framework based on a representative agent model to infer how aggregate selection behavior may vary across new combinatorial choice environments unobserved in historical data. Our first main contribution is an exact characterization of the selection probabilities representable under the s-RAM model over general 0–1 polytopes, and a demonstration that consistency verification can be reduced to solving a polynomial-sized linear program. This characterization, in turn, enables a tractable nonparametric approach for generating predictions in counterfactual environments. For data that are inconsistent with the s-RAM model, we introduce a compact mixed-integer convex formulation to compute the best-fitting s-RAM approximation. We conduct extensive numerical experiments that demonstrate the strong empirical performance of the proposed methods.
We conclude by highlighting several challenges for future research: Developing advanced methods for handling large scale instances where the proposed integer programs formulations may fail to scale well is an important direction for future research. Another direction is to extend the framework to settings where the data is from a polytope for which a 011 representation is not available, such as the Traveling Salesman Problem (TSP), or even beyond 011 polytopes, as in the knapsack problem, thereby accommodating richer combinatorial structures, and broadening the framework’s applicability.
@@ Appendix A Missing Proofs of the Paper
(16) $\nu_{j}-c_{j}^{\prime}(p_{j}^{k})+\sum_{i\in[m]}\lambda_{i}^{k}a_{ij}+\mu_{j}^{k}-w_{j}^{k}=0,\quad\forall j\in\mathcal{I}^{(k)},$
(17) $\mu_{j}^{k}x_{j}^{*}=0,\ w_{j}^{k}(1-x_{j}^{*})=0,\quad\forall j\in\mathcal{I}^{(k)}.$
@@
Figure 6. An illustration of the construction of the functions $c_{j}^{\prime}$ when: (a) $p_{j}^{K_{j}}<1$ and (b) $p_{j}^{1}>0$.
@@
This is possible in polynomial time by solving a linear program where the above conditions are formulated as constraints and maximizing $\epsilon.$ This linear program involves $mK$ variables for $(\lambda_{i}^{k}\in\mathbb{R}:i\in[m],k\in\mathcal{K})$ and one variable for $\epsilon,$ and $\mathcal{O}((n+m)K)$ constraints. ∎
Proof.
@@
Therefore, the optimal value of the formulations in Propositions 4.1 and 4.2 are identical. ∎
Proof.
@@
Therefore, the optimal value of the formulations in Propositions 5.1 and 5.2 are identical. ∎
Appendix B Missing Details on Numerical Experiments in Section 6
For reproducibility, we report the computational environment used in all experiments. Computations were performed on a MacBook Pro with a 12-core Apple M4 Pro chip and 24 GB RAM. The code was executed in Python 3.13.5 with NumPy 2.1.3, SciPy 1.15.3, pandas 2.2.3, and CVXPY 1.7.2. The optimization problems were solved using Gurobi 12.0.3 and MOSEK 11.1.2.
B.1. Data Generation
Gaussian Data for representability and estimation test. We provide a detailed description of the data generation procedure used in our experiments for evaluating the representational power of s-RAM across four combinatorial problems: longest path, shortest path, assignment, and constrained subset selection.
@@ B.1. Data Generation
In the constrained subset selection problem, for each collection of available subsets $\{\mathcal{I}^{(k):k\in\mathcal{K}}\}$, we generate an instance $\boldsymbol{p}^{\mathcal{K}}$ as follows. For each $k\in\mathcal{K}$, the marginal allocation weights $bp^{(k)}$ are generated using the same procedure as in the longest path experiments. Specifically, item costs take the form $v_{i}+\epsilon_{i}$, where the deterministic components $v_{i}$ and noise terms $\epsilon_{i}$ follow the same distributions, noise structures, and parameter settings as before. The only difference is that, for each realization, we solve the corresponding deterministic constrained subset selection problem to obtain the observed marginal allocation weights.
For an instance $\boldsymbol{p}^{\mathcal{K}}$, we verify whether the observed data $\boldsymbol{p}^{\mathcal{K}}$ is representable by s-RAM by checking the conditions in equation (3). If the condition is violated, we compute the average absolute deviation loss using equation (12), where the loss function is defined as $\frac{1}{K}\sum_{k\in\mathcal{K}}\sum_{e\in\boldsymbol{E}^{k}}|x_{e}^{k}-p_{e}^{k}|.$ Here, $p_{e}^{k}$ is the observed flow on edge $e$ in observation $k$, and $x_{e}^{k}$ is the probability estimated under the s-RAM model or for the constrained subset selection problem, $\frac{1}{K}\sum_{k\in\mathcal{K}}\sum_{i\in\mathcal{I}^{(k)}}|x_{i}^{k}-p_{i}^{k}|.$ Here, $p_{i}^{k}$ is the observed marginal weight of item $i$ in observation $k$, and $x_{i}^{k}$ is the probability estimated under the s-RAM model . Because the data are generated randomly, enforcing exact representability can be overly restrictive: instances with vanishingly small loss may be classified as non-representable due to numerical or sampling noise.
Accordingly, we treat an instance as s-RAM representable if its average absolute deviation loss is below $10^{-6}$, and report the representable fraction under this criterion. Finally, we report, for each $K$ and noise setting, the proportion of representable instances and the average absolute deviation loss across the 20 generated instances.
@@
Longest Path Problem. Recall that for the longest path problem, we assume the noise marginals are nonidentical exponential distributions. Given a collection historical graphs $\boldsymbol{G}^{1}=(\boldsymbol{V}^{1},\boldsymbol{E}^{1}),\boldsymbol{G}^{2}=(\boldsymbol{V}^{2},\boldsymbol{E}^{2}),\cdots,\boldsymbol{G}^{K}=(\boldsymbol{V}^{K},\boldsymbol{E}^{K})$, for a graph $\boldsymbol{G}^{k}$ with $k\in\mathcal{K}$, we solve a convex program defined by the corresponding parametric MDM model to obtain the choice probabilities. Let $\boldsymbol{\nu}$ denote the deterministic weight matrix. Let $\lambda_{ij}$ be the rate parameters for each edge $(i,j)$ in the set of all edges $\boldsymbol{E}$. From (Natarajan et al., 2009), we have that, for a graph $\boldsymbol{G}^{k}$, the convex program with nonidentical exponential marginal distributions for the longest path problem is as follows:
@@
(18) $\begin{aligned} \max_{\boldsymbol{x}}&\sum_{(i,j)\in\boldsymbol{E}^{k}}(\nu_{ij}+\frac{1}{\lambda_{ij}})x_{ij}-\sum_{(i,j)\in\boldsymbol{E}^{k}}\frac{1}{\lambda_{ij}}x_{ij}\ln{x_{ij}}\\ \text{s.t.}\;&\sum_{j:(i,j)\in\boldsymbol{E}^{k}}x_{ij}-\sum_{j:(j,i)\in\boldsymbol{E}^{k}}x_{ji}=\begin{cases}1,\quad&i=s,\\ -1,\quad&i=t,\\ 0,\quad&i\in\boldsymbol{V}^{k},\end{cases}\\ &x_{ij}\geq 0\quad\forall\ (i,j)\in\boldsymbol{E}^{k}.\end{aligned}$
These choice probabilities are used as the observed data $\boldsymbol{p}^{\mathcal{K}}$ to train the nonparametric s-RAM model. For prediction, we generate a new graph $\boldsymbol{G}^{\text{new}}$ and compute the probabilities via 18 with $\boldsymbol{G}^{\text{new}}$.
Shortest Path Problem. Recall that for the longest path problem, we assume the noise marginals are nonidentical uniform distributions. Let $\boldsymbol{\nu}$ denote the deterministic weight matrix, and let $\mathbf{a}$ and $\mathbf{b}$ denote the matrices of lower and upper bounds of the uniform noise distributions, respectively. From (Natarajan et al., 2009), we have that, for a graph $\boldsymbol{G}^{k}$ with $k\in\mathcal{K}$, the convex program with nonidentical uniform marginal distributions for the shortest path problem is as follows: